marimartinez:

Pls help me lmao............Which set of ratios could be used to determine if one triangle is a dilation of the other? A triangle has side lengths of 4, 6, 8.5. A second triangle has side lengths of 6, 9, 12.5. StartFraction 4 Over 6 EndFraction = StartFraction 6 Over 9 EndFraction = StartFraction 8.5 Over 12.5 EndFraction StartFraction 6 Over 4 EndFraction = StartFraction 6 Over 9 EndFraction = StartFraction 8.5 Over 12.5 EndFraction StartFraction 4 Over 6 EndFraction = StartFraction 9 Over 6 EndFraction = StartFraction 8.5 Over 12.5 EndFraction StartFraction 4 Over 6 EndFraction = StartFraction 8.5 Over 9 EndFraction = StartFraction 6 Over 12.5 EndFraction

Trickjoker:

Hey! Welcome to QuestionCove, an online learning platform for all ages. I hope you enjoy your stay and feel free to use the site more! Now, moving on to the question... I think it's the StartFractio one.

ScamTheMan:

A simple way to get the ratio of these two triangles would be to put one number of the first triangle over the corresponding number of the other triangle. Then it set it equal and repeat the process. However, the numerators must be from the same triangle, and the denominator would also have to be from the same triangle. We can easily eliminate B and C since the numerators aren't from the same triangle throughout the equation.

ScamTheMan:

Then it's between either: \[4/6 = 6/9 = 8.5/12.5\] (A) and \[4/6=8.5/9=6/12.5\] (D) We can divide through or simplify the fractions to see if they're all the same. If they don't equal each other, then that answer choice isn't correct.

ScamTheMan:

So let's take answer choice D and see if all the fractions are equal to each other. \[4/6=0.67\] \[8.5/9=0.94\] \[6/12.5=0.48\] As you can see the statement: \[0.67=0.95=0.48\] Is clearly not true. Meaning our answer can't be D. Therefore, it's A. And if we were to test it, we would see the numbers are relatively the same.

marimartinez:

Thank you guys💯💯

ScamTheMan:

course <3